The element Chemisfunium (Ch) has an average atomic mass of 91.4381 amu and is composed of two isotopes. The first has a mass of 92.1026 amu and makes up 50.540% of a standard sample. What is the mass of the second isotope?
Answer #1

Sol . As Average atomic mass of element Ch = (Mass of 1st isotope × perecentage abundance of 1st isotope) / 100 + ( Mass of 2nd isotope × percentage abundance of 2nd isotope )/100

Now , percentage abundance of 1st isotope = 50.540 %

percentage abundance of 2nd  isotope = 100 - 50.540

=49.46 %

Mass of 1st isotope = 92.1026 amu

Average atomic mass of element Ch = 91.4381 amu

So , 91.4381 = (92.1026 × 50.540)/100 + (Mass of 2nd isotope × 49.46)/100

91.4381 = 46.54865404 + Mass of 2nd isotope × 0.4946

Therefore , mass of 2nd isotope = (91.4381-46.54865404)/0.4946 =  90.7591 amu